OpenStudy (anonymous):

Could someone please help??? is the graph of f(x)=2(x^1/2)-2A, where A>0, increasing, decreasing, convex, concave, have an inflection pt at x=2A^(-2/3), 4A^(-2/3) or A^(-2/3)? I'm thinking that it's concave, but I was wondering if someone could check it just ot make sure. THanks!

OpenStudy (anonymous):

increasing, solve \[\frac{1}{\sqrt{x}}-A>0\] for \(x\)

OpenStudy (anonymous):

oh that was wrong!!!

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x}}>A\] \[\sqrt{x}<\frac{1}{A}\] \[x<\frac{1}{A^2}\] that is better of course also \(x>0\) but we knew that to begin with

OpenStudy (anonymous):

second derivative has no A in it it is just as before \[f''(x)=-\frac{1}{x^{\frac{3}{2}}}\] and this is always negative

OpenStudy (anonymous):

so always concave down (i guess that is convex)

OpenStudy (anonymous):

oh i see you changed the problem, so maybe my first answer is incorrect i wasn't reading carefully

OpenStudy (anonymous):

huh? I'm really confused, sorry.

OpenStudy (anonymous):

but i think you made a typo, because \[f(x)=2\sqrt{x}-2A\] has exactly the same behaviour as \(f(x)=2\sqrt{x}\) the \(-2A\) makes no difference

OpenStudy (anonymous):

Oh i did, sorry! I meant -Ax

OpenStudy (anonymous):

So i typed this into an online calculator and it as concave down, but i that decreasing because it not the first derivative? or is it concave down? I get confused with that stuff

OpenStudy (anonymous):

ok so we take the first derivative, which we already did

OpenStudy (anonymous):

that is for increasing and decreasing that is what i solved above

OpenStudy (anonymous):

then the second derivative has no A in it, the second derivative is just \[f''(x)=-\frac{1}{x^{\frac{3}{2}}}\]

OpenStudy (anonymous):

the A is gone, so the concavity is independent of A the second derivative is negative for all values of \(x\) therefore the function is always concave down

OpenStudy (anonymous):

ok, thanks! that makes sense

OpenStudy (anonymous):

I have one more question. Find f'(1) if f(x)=the integral (1+(X^3))^(1/2). Is this a trick question? cause it's asking for the derivative, and the integral is already in derivative form, would i just plug in 1 and get the square root of 2?